The meaning of LIGHT (and E&M!): MAXWELL’S EQUATION

▽ · D = ρ , ▽ · B = 0 , ▽ × E = −\frac{∂B}{∂t} , ▽ × H = J + \frac{∂D}{∂t} .

E: electric field; H: magnetic field ; D: electric displacement field ; B: magnetic flux density ; ρ: electric charge density; J: free current density.